# A class of forward-backward regularizations of the Perona-Malik equation with variable exponent

### v1

@@ 2.2 Young measures
Lemma 2.8.
[47] Let $\Omega\subset\mathbb{R}^{N}$ be Lebesgue measurable, let $K\subset\mathbb{R}^{N}$ be closed, and let $u_{j}:\Omega\rightarrow\mathbb{R}^{N},j\in\mathbb{N}$, be a sequence of Lebesgue measurable functions satisfying $u_{j}\rightarrow K$ in measure as $j\rightarrow\infty$, i.e., given any open neighbourhood $U$ of $K$ in $\mathbb{R}^{N}$
@@
Definition 2.3.
The map $\nu:\Omega\rightarrow\mathcal{M}(\mathbb{R}^{N})$ in Lemma 2.8 is called the Young measure in $\mathbb{R}^{N}$ generated by the sequence $\{u_{j_{k}}\}_{k=1}^{\infty}$.
Hungerbühler subsequently established the following consequence, while refined versions of Lemma 2.8 can be found in [62].
Lemma 2.9.
[63] Let $\Omega$, $u_{j}$, and $\nu_{x}$ be as in Lemma 2.8. Then (15), (16) and (17) are equivalent.
Remark 1.
In [47], it is also shown that, under hypothesis (15), for any measurable set $D\subset\Omega$,
@@ 2.3 Definition of Young measure solutions
Our goal is to construct a Young measure solution to problem (1)–(3). Following the approach in [26, 25], we first consider the homogeneous Neumann IBVP for the Sobolev regularization of equation (1):
@@ 3 Weak solutions of the regularized problem (22)–(24)
which shows that the mapping $\xi\mapsto\varphi_{\varepsilon}(x,\xi)$ is convex for every $x\in\Omega$ whenever $m\geq\frac{KT}{\varepsilon}$. Lemma 2.7 then yields the weak lower semicontinuity of $u\mapsto\int_{\Omega}\varphi_{\varepsilon}(x,\nabla u)dx$ on $W^{1,p(\cdot)}(\Omega)$. Let $\{u_{m}^{j,k}\}_{k=1}^{\infty}\subset W_{f}^{1,p(\cdot)}(\Omega)$ be a minimizing sequence of $J^{m,j}$. Since $J^{m,j}$ is coercive on $W_{f}^{1,p(\cdot)}(\Omega)$ by Lemma 2.5, we may assume that $\{u_{m}^{j,k}\}_{k=1}^{\infty}$ is bounded in $W^{1,p(\cdot)}(\Omega)$. Applying Lemma 2.2, there exists $u_{m}^{j}\in W_{f}^{1,p(\cdot)}(\Omega)$ and a subsequence of $\{u_{m}^{j,k}\}_{k=1}^{\infty}$, still denoted by itself, such that $u_{m}^{j,k}\rightharpoonup u_{m}^{j}$ weakly in $W^{1,p(\cdot)}(\Omega)$. Then, we have
@@
Consequently, $u_{m}^{j}\in W_{f}^{1,p(\cdot)}(\Omega)$ is a (not necessarily unique) minimizer of $J^{m,j}$. This finishes the proof. ∎
@@
From (18) and the facts that $\nabla u_{m}^{j}\in L^{p(\cdot)}(\Omega)$ and $\vec{q}(x,\nabla u_{m}^{j})\in L^{p^{\prime}(\cdot)}(\Omega)$, together with Lemma 2.4, it follows that $\phi\in W^{1,p(\cdot)}(\Omega)$ is admissible as a test function in the the Euler-Lagrange system (28) for $u_{m}^{j}$. We now derive an a priori estimate for the discrete energy.
Lemma 3.3 (Discrete energy inequality).
@@
$\int_{\Omega}\varphi(x,\nabla u_{m}^{j})dx+$ $\sum_{i=1}^{j}\int_{\Omega}\frac{m}{T}\left|u_{m}^{i}-u_{m}^{i-1}\right|^{2}dx$
$+$ $\sum_{i=1}^{j}\int_{\Omega}(\varepsilon-\gamma)\frac{m}{T}\left|\nabla u_{m}^{i}-\nabla u_{m}^{i-1}\right|^{2}dx\leq\int_{\Omega}\varphi(x,\nabla f)dx,$ (29)
@@
Set $m_{0}=\max\left\{\frac{KT}{2\gamma},\frac{KT}{\varepsilon}\right\}$. Then, by Lemma 3.2, $J^{m,j}$ admits a minimizer $u_{m}^{j}\in W_{f}^{1,p(\cdot)}(\Omega)$, satisfying (28) for every $m\geq m_{0}$ and $j=1,2,\cdots,m$. From (30), for $i=1,2,\cdots,j$, we derive
@@
Before proceeding with the estimate, we consider equation (28) with $\phi=u_{m}^{i}-u_{m}^{i-1}$, which yields the following expression
@@
$\leq$ $-\int_{\Omega}\frac{m}{T}\left|u_{m}^{i}-u_{m}^{i-1}\right|^{2}dx-\int_{\Omega}(\varepsilon-\gamma)\frac{m}{T}\left|\nabla u_{m}^{i}-\nabla u_{m}^{i-1}\right|^{2}dx.$
@@
$\int_{\Omega}\varphi(x,\nabla u_{m}^{j})dx-\int_{\Omega}\varphi(x,\nabla f)dx$
$\leq$ $-\sum_{i=1}^{j}\int_{\Omega}\frac{m}{T}\left|u_{m}^{i}-u_{m}^{i-1}\right|^{2}dx-\sum_{i=1}^{j}\int_{\Omega}(\varepsilon-\gamma)\frac{m}{T}\left|\nabla u_{m}^{i}-\nabla u_{m}^{i-1}\right|^{2}dx.$
This completes the proof. ∎
Corollary 3.4.
@@
$\sup_{1\leq j\leq m}\left\|u_{m}^{j}\right\|_{W^{1,p(\cdot)}(\Omega)}+$ $\sum_{j=1}^{m}\frac{m}{T}\left\|u_{m}^{j}-u_{m}^{j-1}\right\|_{L^{2}(\Omega)}^{2}$
$+$ $\sum_{j=1}^{m}\left(\varepsilon-\gamma\right)\frac{m}{T}\left\|\nabla u_{m}^{j}-\nabla u_{m}^{j-1}\right\|_{L^{2}(\Omega)}^{2}\leq M_{1},$ (32)
@@
The assertion follows directly from the growth condition (18), Lemma 2.5 and 3.3, and the assumption that $f\in W^{1,p(\cdot)}(\Omega)$.
@@
for every $\zeta\in L^{1}(0,T;W^{1,p(\cdot)}(\Omega))\cap L^{2}(0,T;H^{1}(\Omega))$. To establish the existence of weak solutions to problem (22)–(24), we use Corollary 3.4 to obtain uniform estimates for the approximate solutions.
Lemma 3.5.
@@
and, analogously, we obtain the corresponding estimate for $\bar{u}_{m}$. Furthermore, we derive from (18) and Lemma 2.1 that
@@
$\leq$ $2^{\frac{1}{p^{-}-1}}\left(\lambda_{2}+1\right)^{\frac{p^{-}}{p^{-}-1}}\max\left\{\left\|\nabla u_{m}^{j}\right\|_{L^{p(\cdot)}(\Omega)}^{p^{-}},\left\|\nabla u_{m}^{j}\right\|_{L^{p(\cdot)}(\Omega)}^{p^{+}}\right\}$
$+2^{\frac{1}{p^{-}-1}}\left|\Omega\right|$
@@
The proof is thus complete. ∎
By Lemma 2.2, upon extracting suitable (not relabeled) subsequences, we obtain the following convergences as $m\to\infty$:
@@
For any $t\in[0,T)$, there exists an index $l\in\{1,2,\cdots,m\}$ such that $(l-1)\frac{T}{m}\leq t<l\frac{T}{m}$. For the case $l\geq 2$, from (32), (40), and the Hölder inequality (see Lemma 2.3), we derive
$\int_{0}^{t}\int_{\Omega}\left|\nabla\bar{u}_{m}-\nabla u\right|^{2}dxd\tau$
$=$ $\sum_{i=1}^{l-1}\int_{(i-1)\frac{T}{m}}^{i\frac{T}{m}}\int_{\Omega}\left|\nabla u_{m}^{i}-\nabla u\right|^{2}dxdt+\int_{(l-1)\frac{T}{m}}^{t}\int_{\Omega}\left|\nabla\bar{u}_{m}-\nabla u\right|^{2}dxd\tau$
@@
We denote the three terms on the left-hand side of the above equality by $T_{1}$, $T_{2}$, and $T_{3}$, and write $\theta(m)$ for quantities tending to zero as $m\rightarrow\infty$, uniformly with respect to $t$. By (19) and Lemma 2.6 (e.g., with $\varPsi=\{\chi_{Q_{t}}\vec{q}(x,\nabla u):t\in[0,T]\}$), we obtain
@@
The proof of Lemma 3.6 is complete. ∎
@@
$=$ $\lim_{m\rightarrow\infty}\bigg(-\iint_{Q_{T}}\frac{\partial u_{m}}{\partial t}\bar{u}_{m}dxdt-\iint_{Q_{T}}\frac{\left|\nabla\bar{u}_{m}\right|^{2}}{1+\left|\nabla\bar{u}_{m}\right|^{2}}dxdt$
$\quad-\iint_{Q_{T}}\varepsilon\nabla\frac{\partial u_{m}}{\partial t}\cdot\nabla\bar{u}_{m}dxdt\bigg)$
@@
Thus, applying the standard monotone method, we obtain $\vec{\sigma}=\left|\nabla u\right|^{p(x)-2}\nabla u$ a.e. in $Q_{T}$, which verifies that (45) holds. ∎
@@
To prove Theorem 3.1, it remains to establish the energy estimate (26). By Lemma 3.3, for any $\gamma\in(0,\varepsilon)$, there exists $m_{0}(\gamma)>0$ such that, for all $m\geq m_{0}$ and for a.e. $t\in[0,T)$, we derive from (29) the following inequalities, where $(j-1)\frac{T}{m}\leq t<j\frac{T}{m}$ holds for some $j$:
@@
From Lemma 3.6, we know that $\nabla\bar{u}_{m}(t)\rightarrow\nabla u(t)$ a.e. in $\Omega$. Then, Fatou’s Lemma implies that
@@
for a.e. $t\in[0,T]$. This completes the proof of Theorem 3.1.
4 Proof of Theorem 1.1
In this section, we prove the existence of Young measure solutions to problem (1)–(3) using Theorem 3.1 and the corresponding limit process $\varepsilon\to 0^{+}$.
For any $\varepsilon\in(0,1)$, it follows from Theorem 3.1 that the regularized problem (22)–(24) admits a unique weak solution $u^{\varepsilon}\in L^{\infty}(0,T;W_{f}^{1,p(\cdot)}(\Omega))\cap H^{1}(0,T;H^{1}(\Omega))$, which satisfies
@@
By Lemmas 2.1 and 2.5, along with the mean invariance of $u^{\varepsilon}(t)$, we deduce that
$\left\|u^{\varepsilon}_{t}\right\|_{L^{2}(Q_{T})}+\left\|\nabla\sqrt{\varepsilon}u^{\varepsilon}_{t}\right\|_{L^{2}(Q_{T})}+$ $\sup_{0\leq t\leq T}\left\|u^{\varepsilon}(t)\right\|_{W^{1,p(\cdot)}(\Omega)}$
$+$ $\sup_{0\leq t\leq T}\left\|\vec{q}\left(x,\nabla u^{\varepsilon}(t)\right)\right\|_{L^{p^{\prime}(\cdot)}(\Omega)}\leq M,$ (52)
@@
For a.e. $t\in[0,T]$, let $\nu_{\cdot,t}=(\nu_{x,t})_{x\in\Omega}$ denote the Young measure generated by the sequence $\{\nabla u^{\varepsilon}(t)\}_{0<\varepsilon<1}$. Since the generating sequence satisfies the boundedness condition (15), it follows from Lemmas 2.8 and 2.9 that $\left\|\nu_{x,t}\right\|_{\mathcal{M}(\mathbb{R}^{N})}=1$ a.e. in $\Omega$. We now verify that the pair $\left(u,\nu\right)$ is a Young measure solution to problem (1)–(3). Combining Remark 1 with (52), we infer that
@@
for every $\zeta\in L^{1}(0,T;W^{1,p(\cdot)}(\Omega))\cap L^{2}(0,T;H^{1}(\Omega))$. Moreover, $u(0)=f$ a.e. in $\Omega$ follows from the same argument as in the proof of Lemma 3.6, and the details are omitted here. Therefore, $\left(u,\nu\right)$ is a Young measure solution to problem (1)–(3), and the proof is complete.
[…]
Acknowledgements The authors wish to express their sincere gratitude to Lecturer Jingfeng Shao (Guangxi University) for many helpful discussions during the preparation of this paper. They also thank the reviewers and editors for their valuable and insightful comments.
[…]
Funding Z. Guo is supported by NSFC (U21B2075, 12171123) and Fundamental Research Funds for the Central Universities (2022FRFK060020). W. Yao is supported by NSFC (12301536) and Fundamental Research Funds for the Central Universities (HIT.NSRIF202302).
[…]
Data Availability Data sharing is not applicable to this article as no datasets were generated or analysed during the current study.
Declarations
[…]
Conflict of interest The authors declared no potential conflict of interest with respect to the research, authorship, and publication of this article.
[…]
Ethics approval The research does not involve humans and/or animals. The authors declare that there are no ethics issues to be approved or disclosed.

### v2

@@ 2.2 Young measures
Lemma 8
[ball2005version] Let $\Omega\subset\mathbb{R}^{N}$ be Lebesgue measurable, let $K\subset\mathbb{R}^{N}$ be closed, and let $u_{j}:\Omega\rightarrow\mathbb{R}^{N},j\in\mathbb{N}$, be a sequence of Lebesgue measurable functions satisfying $u_{j}\rightarrow K$ in measure as $j\rightarrow\infty$, i.e., given any open neighbourhood $U$ of $K$ in $\mathbb{R}^{N}$
@@
Definition 3
The map $\nu:\Omega\rightarrow\mathcal{M}(\mathbb{R}^{N})$ in Lemma 8 is called the Young measure in $\mathbb{R}^{N}$ generated by the sequence $\{u_{j_{k}}\}_{k=1}^{\infty}$.
Hungerbühler subsequently established the following consequence, while refined versions of Lemma 8 can be found in [Kinderlehrer1994059].
Lemma 9
[Hungerbuhler1997] Let $\Omega$, $u_{j}$, and $\nu_{x}$ be as in Lemma 8. Then (14), (15) and (16) are equivalent.
Remark 1
In [ball2005version], it is also shown that, under hypothesis (14), for any measurable set $D\subset\Omega$,
@@ 2.3 Definition of Young measure solutions
Our goal is to construct a Young measure solution to problem (1)–(3). Following the approach in [Oliver20031380, THANH20141403], we first consider the homogeneous Neumann IBVP for the Sobolev regularization of equation (1):
@@ 3 Weak solutions of the regularized problem (21)–(23)
which shows that the mapping $\xi\mapsto\varphi_{\varepsilon}(x,\xi)$ is convex for every $x\in\Omega$ whenever $m\geq\frac{KT}{\varepsilon}$. Lemma 7 then yields the weak lower semicontinuity of $u\mapsto\int_{\Omega}\varphi_{\varepsilon}(x,\nabla u)dx$ on $W^{1,p(\cdot)}(\Omega)$. Let $\{u_{m}^{j,k}\}_{k=1}^{\infty}\subset W_{f}^{1,p(\cdot)}(\Omega)$ be a minimizing sequence of $J^{m,j}$. Since $J^{m,j}$ is coercive on $W_{f}^{1,p(\cdot)}(\Omega)$ by Lemma 5, we may assume that $\{u_{m}^{j,k}\}_{k=1}^{\infty}$ is bounded in $W^{1,p(\cdot)}(\Omega)$. Applying Lemma 2, there exists $u_{m}^{j}\in W_{f}^{1,p(\cdot)}(\Omega)$ and a subsequence of $\{u_{m}^{j,k}\}_{k=1}^{\infty}$, still denoted by itself, such that $u_{m}^{j,k}\rightharpoonup u_{m}^{j}$ weakly in $W^{1,p(\cdot)}(\Omega)$. Then, we have
@@
Consequently, $u_{m}^{j}\in W_{f}^{1,p(\cdot)}(\Omega)$ is a (not necessarily unique) minimizer of $J^{m,j}$. This finishes the proof.
@@
From (17) and the facts that $\nabla u_{m}^{j}\in L^{p(\cdot)}(\Omega)$ and $\vec{q}(x,\nabla u_{m}^{j})\in L^{p^{\prime}(\cdot)}(\Omega)$, together with Lemma 4, it follows that $\phi\in W^{1,p(\cdot)}(\Omega)$ is admissible as a test function in the the Euler-Lagrange system (27) for $u_{m}^{j}$. We now derive an a priori estimate for the discrete energy.
Lemma 11 (Discrete energy inequality)
@@
$\int_{\Omega}\varphi(x,\nabla u_{m}^{j})dx+\sum_{i=1}^{j}\int_{\Omega}\frac{m}{T}\left|u_{m}^{i}-u_{m}^{i-1}\right|^{2}dx+\sum_{i=1}^{j}\int_{\Omega}(\varepsilon-\gamma)\frac{m}{T}\left|\nabla u_{m}^{i}-\nabla u_{m}^{i-1}\right|^{2}dx\leq\int_{\Omega}\varphi(x,\nabla f)dx,$ (28)
@@
Set $m_{0}=\max\left\{\frac{KT}{2\gamma},\frac{KT}{\varepsilon}\right\}$. Then, by Lemma 10, $J^{m,j}$ admits a minimizer $u_{m}^{j}\in W_{f}^{1,p(\cdot)}(\Omega)$, satisfying (27) for every $m\geq m_{0}$ and $j=1,2,\cdots,m$. From (29), for $i=1,2,\cdots,j$, we derive
@@
Before proceeding with the estimate, we consider (27) with $\phi=u_{m}^{i}-u_{m}^{i-1}$, which yields
@@
$\int_{\Omega}\varphi(x,\nabla u_{m}^{i})dx-\int_{\Omega}\varphi(x,\nabla u_{m}^{i-1})dx\leq-\int_{\Omega}\frac{m}{T}\left|u_{m}^{i}-u_{m}^{i-1}\right|^{2}dx-\int_{\Omega}(\varepsilon-\gamma)\frac{m}{T}\left|\nabla u_{m}^{i}-\nabla u_{m}^{i-1}\right|^{2}dx.$
@@
$\int_{\Omega}\varphi(x,\nabla u_{m}^{j})dx-\int_{\Omega}\varphi(x,\nabla f)dx\leq-\sum_{i=1}^{j}\int_{\Omega}\frac{m}{T}\left|u_{m}^{i}-u_{m}^{i-1}\right|^{2}dx-\sum_{i=1}^{j}\int_{\Omega}(\varepsilon-\gamma)\frac{m}{T}\left|\nabla u_{m}^{i}-\nabla u_{m}^{i-1}\right|^{2}dx.$
This completes the proof.
Corollary 1
@@
$\sup_{1\leq j\leq m}\left\|u_{m}^{j}\right\|_{W^{1,p(\cdot)}(\Omega)}+\sum_{j=1}^{m}\frac{m}{T}\left\|u_{m}^{j}-u_{m}^{j-1}\right\|_{L^{2}(\Omega)}^{2}+\sum_{j=1}^{m}\left(\varepsilon-\gamma\right)\frac{m}{T}\left\|\nabla u_{m}^{j}-\nabla u_{m}^{j-1}\right\|_{L^{2}(\Omega)}^{2}\leq M_{1},$ (31)
@@
The assertion follows directly from the growth condition (17), Lemma 5 and 11, and the assumption that $f\in W^{1,p(\cdot)}(\Omega)$.
@@
for every $\zeta\in L^{1}(0,T;W^{1,p(\cdot)}(\Omega))\cap L^{2}(0,T;H^{1}(\Omega))$. To establish the existence of weak solutions to problem (21)–(23), we use Corollary 1 to obtain uniform estimates for the approximate solutions.
Lemma 12
@@
and, analogously, we obtain the corresponding estimate for $\bar{u}_{m}$. Furthermore, we derive from (17) and Lemma 1 that
@@
$\leq$ $2^{\frac{1}{p^{-}-1}}\left(\lambda_{2}+1\right)^{\frac{p^{-}}{p^{-}-1}}\max\left\{\left\|\nabla u_{m}^{j}\right\|_{L^{p(\cdot)}(\Omega)}^{p^{-}},\left\|\nabla u_{m}^{j}\right\|_{L^{p(\cdot)}(\Omega)}^{p^{+}}\right\}+2^{\frac{1}{p^{-}-1}}\left|\Omega\right|$
@@
The proof is thus complete.
By Lemma 2, upon extracting suitable (not relabeled) subsequences, we obtain the following convergences as $m\to\infty$:
@@
For any $t\in[0,T)$, there exists an index $l\in\{1,2,\cdots,m\}$ such that $(l-1)\frac{T}{m}\leq t<l\frac{T}{m}$. For the case $l\geq 2$, from (31), (39), and the Hölder inequality (see Lemma 3), we derive
$\int_{0}^{t}\int_{\Omega}\left|\nabla\bar{u}_{m}-\nabla u\right|^{2}dxd\tau=$ $\sum_{i=1}^{l-1}\int_{(i-1)\frac{T}{m}}^{i\frac{T}{m}}\int_{\Omega}\left|\nabla u_{m}^{i}-\nabla u\right|^{2}dxdt+\int_{(l-1)\frac{T}{m}}^{t}\int_{\Omega}\left|\nabla\bar{u}_{m}-\nabla u\right|^{2}dxd\tau$
@@
We denote the three terms on the left-hand side of the above equality by $T_{1}$, $T_{2}$, and $T_{3}$, and write $\theta(m)$ for quantities tending to zero as $m\rightarrow\infty$, uniformly with respect to $t$. By (18) and Lemma 6 (e.g., with $\varPsi=\{\chi_{Q_{t}}\vec{q}(x,\nabla u):t\in[0,T]\}$), we obtain
@@
The proof of Lemma 13 is complete.
@@
$=$ $\lim_{m\rightarrow\infty}\left(-\iint_{Q_{T}}\frac{\partial u_{m}}{\partial t}\bar{u}_{m}dxdt-\iint_{Q_{T}}\frac{\left|\nabla\bar{u}_{m}\right|^{2}}{1+\left|\nabla\bar{u}_{m}\right|^{2}}dxdt-\iint_{Q_{T}}\varepsilon\nabla\frac{\partial u_{m}}{\partial t}\cdot\nabla\bar{u}_{m}dxdt\right)$
@@
Thus, applying the standard monotone method, we obtain $\vec{\sigma}=\left|\nabla u\right|^{p(x)-2}\nabla u$ a.e. in $Q_{T}$, which verifies that (44) holds.
@@
To prove Theorem 2, it remains to establish the energy estimate (25). By Lemma 11, for any $\gamma\in(0,\varepsilon)$, there exists $m_{0}(\gamma)>0$ such that, for all $m\geq m_{0}$ and for a.e. $t\in[0,T)$, we derive from (28) the following inequalities, where $(j-1)\frac{T}{m}\leq t<j\frac{T}{m}$ holds for some $j$:
@@
From Lemma 13, we know that $\nabla\bar{u}_{m}(t)\rightarrow\nabla u(t)$ a.e. in $\Omega$. Then, Fatou’s Lemma implies that
@@
for a.e. $t\in[0,T]$. This completes the proof of Theorem 2.
4 Proof of Theorem 1
In this section, we prove the existence of Young measure solutions to problem (1)–(3) using Theorem 2 and the corresponding limit process $\varepsilon\to 0^{+}$.
For any $\varepsilon\in(0,1)$, it follows from Theorem 2 that the regularized problem (21)–(23) admits a unique weak solution $u^{\varepsilon}\in L^{\infty}(0,T;W_{f}^{1,p(\cdot)}(\Omega))\cap H^{1}(0,T;H^{1}(\Omega))$, which satisfies
@@ 4 Proof of Theorem 1
By Lemmas 1 and 5, along with the mean invariance of $u^{\varepsilon}(t)$, we deduce that
$\left\|u^{\varepsilon}_{t}\right\|_{L^{2}(Q_{T})}+\left\|\nabla\sqrt{\varepsilon}u^{\varepsilon}_{t}\right\|_{L^{2}(Q_{T})}+\sup_{0\leq t\leq T}\left\|u^{\varepsilon}(t)\right\|_{W^{1,p(\cdot)}(\Omega)}+\sup_{0\leq t\leq T}\left\|\vec{q}\left(x,\nabla u^{\varepsilon}(t)\right)\right\|_{L^{p^{\prime}(\cdot)}(\Omega)}\leq M,$ (51)
@@
For a.e. $t\in[0,T]$, let $\nu_{\cdot,t}=(\nu_{x,t})_{x\in\Omega}$ denote the Young measure generated by the sequence $\{\nabla u^{\varepsilon}(t)\}_{0<\varepsilon<1}$. Since the generating sequence satisfies the boundedness condition (14), it follows from Lemmas 8 and 9 that $\left\|\nu_{x,t}\right\|_{\mathcal{M}(\mathbb{R}^{N})}=1$ a.e. in $\Omega$. We now verify that the pair $\left(u,\nu\right)$ is a Young measure solution to problem (1)–(3). Combining Remark 1 with (51), we infer that
@@
for every $\zeta\in L^{1}(0,T;W^{1,p(\cdot)}(\Omega))\cap L^{2}(0,T;H^{1}(\Omega))$. Moreover, $u(0)=f$ a.e. in $\Omega$ follows from the same argument as in the proof of Lemma 13, and the details are omitted here. Therefore, $\left(u,\nu\right)$ is a Young measure solution to problem (1)–(3), and the proof is complete.
Acknowledgements
The authors wish to express their sincere gratitude to Lecturer Jingfeng Shao (Guangxi University) for many helpful discussions during the preparation of this paper. This work was supported by the National Natural Science Foundation of China (Grant Nos. U21B2075, 12171123, 12301536) and the Fundamental Research Funds for the Central Universities (Grant Nos. 2022FRFK060020, HIT.NSRIF202302).
Data availability
No data was used for the research described in the article.